3. X Company is evaluating its cost of capital under alternative financing arrangements. In consultation with investment bankers, X Company expects to be able to issue new debt at par with a coupon rate of 10% and to issue new preferred stock with a $4.00 per share dividend at $25 a share. The common stock of X Company is currently selling for $20.00 a share. X Company expects to pay a dividend of $2.50 per share next year. Market analysts foresee a growth in dividends in Invest stock at a rate of 5% per year. Invest does not expect its cost of debt, preferred stock or common stock, to be different under the two possible financing arrangements. X Company marginal tax rate is 40%. Hint: coupon rate of the bond is the same as the before-tax cost of debt. The two arrangements are: Financing Arrangement, Debt, Preferred Stock, Common Stock, respectively

1, 20% 30% 50%

2, 50% 30% 20%

A. What is the weighted average cost of capital to X Company under the first financing arrangement?


1
Expert's answer
2022-01-16T13:33:19-0500

Let the bond rate be 10%. the share of preferred shares is 20%, ordinary shares 50%, debt 30%:

"WACC=0.2\times\frac{4}{25}\times100+0.5(\frac{2.5}{20}\times100+5)+0.3\times10(1-0.4)=13.75"